A few years ago, a total of 2693 thousand people lived in the metropolitan areas of Las Vegas, Nevada, and Sacremento, California. Sacremento had 285 thousand more residents than Las Vegas. What was the population of each metropolitan area?
What was the population of the Sacremento area?
___ thousand
What was the population of the Las Vegas area?
___ thousand

Answer by macston(5194)

k=1000
.
L=population of Las Vegas; S=population of Sacremento=L+285k
.
L+S=2693k
L+(L+285k)=2693k
2L=2408k
L=1204
ANSWER: Population of Las Vegas was 1204 thousand.
.
S+L+285k=1204k+285k=1489k
ANSWER: Population of Sacremento was 1489 thousand.